Johnathon's household frequently purchases tomatoes from the local market. His demand for tomatoes is:

Qd^J (p)=15-3p

A) When Jonathan went to the market last week the price of 1 tomato was $6. This week the price of 1 tomato is $2.

What is Johnathans original and new consumer surplus?

What is the change in consumer surplus?

B) Johnathan lives with his sister, who also ears tomatoes. She has the following demand function:

Qd^s (p)=24-3p

If Jonathan and his sister are the only members of the household who eat tomatoes what is the household demander tomatoes? (Same as market demand)

C) Graph the household demand for tomatoes.
Answer #1

Jo mathans demand : Qat= 15-3p MD Ini tidly price of bonito. (46=p). Od = 15-3 p = 15-36]=-B.co. So, Janethan ako.will not pGraph of household demand Q = 39-6p. At ipao Q=39 At Q=0; p= 34 13 = 6.5 P= 6.5 . 2 Price a y Cuantity of tomita